OpenStudy (anonymous):

(12x^2 + 9x – 7) ÷ 3x Help, I don't understand.

OpenStudy (anonymous):

Wow i need help on the same exact question...

OpenStudy (espex):

\[\frac{12x^{2}+9x-7}{3x}\rightarrow \frac{12x^{2}}{3x}+\frac{9x}{3x}-\frac{7}{3x}\] and simplify.

OpenStudy (anonymous):

Thank you, though I'm still not sure I understand.

OpenStudy (espex):

What is it that you are asked to do with this question?

OpenStudy (anonymous):

It simply says to multiply or divide. This is division, though I'm not understand how you got to your conclusion or how to reduce further.

OpenStudy (espex):

Are you familiar with fractions?

OpenStudy (anonymous):

Somewhat, I haven't had to do them for a few years now. I think to simplify you see what factors the numerator and denominator have until you reach the point where it can't reduce anymore.

OpenStudy (espex):

Basically, yes.

OpenStudy (espex):

So how many times can 3x go into 12x^2?

OpenStudy (anonymous):

Thank You @eSpeX

OpenStudy (espex):

Umm. You're welcome @Justin_Bieber1994 :D

OpenStudy (anonymous):

3 goes into 12 four times, though I don't know what to do with the square. or erm. ^2 that.

OpenStudy (espex):

What happens if you multiply x*x?

OpenStudy (anonymous):

You either get 0 or 1..I forget..

OpenStudy (espex):

\[2*2=4,\rightarrow 4=2^2\] The same thing applies with 'x'

OpenStudy (espex):

So,\[3x*4x=12x^2\]

OpenStudy (anonymous):

Okay, I'm still unsure as of how to apply that to answer my equation...

OpenStudy (espex):

What do you get when you divide your 12x^2 by 3x?

OpenStudy (anonymous):

4x

OpenStudy (espex):

Exactly, you have just reduced the first term. :)

OpenStudy (espex):

How about the second one \[\frac{9x}{3x}\]

OpenStudy (anonymous):

3x, but does the ^2 carry over with the four?

OpenStudy (espex):

No, you treat each term individually. You are done with the first term, it reduces to 4x, now just worry about the second term. What times 3x = 9x?

OpenStudy (anonymous):

3x

OpenStudy (espex):

3x*3x=9x^2

OpenStudy (anonymous):

3?

OpenStudy (espex):

Very good, now you have the first two terms completed. What can you do with the last term?

OpenStudy (anonymous):

I don't know. 7 doesn't go into three and three doesn't go into 7.

OpenStudy (espex):

Exactly, so there is nothing you can do with that. So your resulting expression is what?

OpenStudy (anonymous):

4x, 3 and 7/3x